### v2

# StorySparkQA: Expert-Annotated QA Pairs with Real-World Knowledge for Children’s Story-Based Learning

Abstract
Interactive story reading is a common parent-child activity, where parents expect to teach both language skills and real-world knowledge beyond the story. While increasing storytelling and reading systems have been developed for this activity, they often fail to infuse real-world knowledge into the conversation. This limitation can be attributed to the existing question-answering (QA) datasets used for children’s education, upon which the systems are built, failing to capture the nuances of how education experts think when conducting interactive story reading activities. To bridge this gap, we design an annotation framework, empowered by existing knowledge graph to capture experts’ annotations and thinking process, and leverage this framework to construct StorySparkQA dataset, which comprises $5,868$ expert-annotated QA pairs with real-world knowledge. We conduct automated and human expert evaluations across various QA pair generation settings to demonstrate that our StorySparkQA can effectively support models in generating QA pairs that target real-world knowledge beyond story content. StorySparkQA  [^1] The StorySparkQA dataset and annotation framework can also be found in: https://github.com/neuhai/StorySparkQA is available at https://huggingface.co/datasets/NEU-HAI/StorySparkQA.
[…]
booktabs
StorySparkQA: Expert-Annotated QA Pairs with Real-World Knowledge for Children’s Story-Based Learning
Jiaju Chen East China Normal University                        Yuxuan Lu Northeastern University                        Shao Zhangthanks: Work done when visiting Northeastern University. Shanghai Jiao Tong University
Bingsheng Yao Northeastern University                        Yuanzhe Dong Stanford University                        Ying Xu Harvard University
Yunyao Li Adobe                        Qianwen Wang University of Minnesota Twin Cities                        Dakuo Wang Northeastern University
Yuling Sunthanks:  Corresponding Author: ylsun@cs.ecnu.edu.cn. East China Normal University
1 Introduction
Figure 1: An example of StorySparkQA dataset. In each story section, educational experts select a concept word, link it to a desired external real-world knowledge, and write an appropriate QA pair. Additional data examples of StorySparkQA are presented in Appendix A.1.
Interactive story reading is a common parent-child activity, where parents often sit together with preschool children, read storybooks, and proactively engage in question-answering (QA) conversations with them Wright (1995); Isbell et al. (2004). Such guided conversations are typically grounded in but beyond the story narratives Kotaman (2013), with parents’ expectations of guiding children to learn real-world knowledge and improving their historical, cultural, and emotional awareness Sun et al. (2024). This immersive story-based interaction has been proven to be effective in better supporting preschooler knowledge learning Zhang et al. (2024), enhancing their reading comprehension capabilities Xu et al. (2021), etc.
Despite the benefits of interactive story reading, parents often struggle to appropriately conduct such interactive story reading with children because of multi-facet difficulties Golinkoff et al. (2019); Sun et al. (2024). Specifically, such interactive story reading needs parents to identify the knowledge of interest during storytelling, formulate the real-world knowledge piece they want to teach in mind ("what to ask"), then ask an engaging question ("how to ask") to children at the appropriate time ("when to ask"). Yet, most parents lack the educational expertise necessary to guide such educational conversations Golinkoff et al. (2019); Sun et al. (2024). Also, parents in contemporary society often hardly maintain high concentration to accompany their children due to the need to deal with other work and family chores at the same time Zhang et al. (2022); Sun et al. (2024).
Recently, AI-assisted storytelling systems (e.g. StoryBuddy Zhang et al. (2022), TaleMate Vargas-Diaz et al. (2023), MatheMyths Zhang et al. (2024)), have demonstrated effectiveness in children’s storytelling scenarios Dietz et al. (2021). These systems primarily take advantage of the verbal communication interface and advanced language models to support natural conversation with humans Mahmood et al. (2023); Chan et al. (2024); Yang et al. (2024); Dietz et al. (2021). Nevertheless, existing AI-assisted storytelling and reading systems are not without limitations. Particularly, building on top of data resources with mostly extractive QA pairs (e.g., FairytaleQA Xu et al. (2022)) – where the answers can be found directly in the story narrative – these systems fall short in helping parents teach real-world knowledge beyond the story narrative Yao et al. (2021), which is one of the main expectations of parents Sun et al. (2024).
We believe a promising approach to bridge this gap is to effectively and exhaustively collect education experts’ knowledge, including their step-by-step thinking process as well as the appropriate QA pairs as final artifacts. Nevertheless, to the best of our knowledge, no such data resources exist in the domain of children education. Further, the collection of such data resources requires annotators to recall a comprehensive and systematic external knowledge range for a given story text, which is challenging even for education experts Berry et al. (2016). As a result, this work aims to facilitate experts’ large-coverage knowledge collection and data annotation, and build an expert-labeled, large-scale QA dataset to support story-based educational QA generation with tri-fold contributions:
  • We designed an annotation framework empowered by ConceptNetSpeer et al. (2017), a knowledge graph (KG) of structured real-world knowledge, to facilitate education experts creating appropriate story-based educational QA pairs, while collecting experts’ mental procedures during data annotation.
  • Based on the proposed annotation framework, we build StorySparkQA, an expert-labeled QA dataset consisting of $5,868$ story-based QA pairs infused with real-world knowledge.
  • We demonstrate the utility of our StorySparkQA on the QA pair generation (QAG) task, benchmarked with a set of popular language models (fine-tuned T5-Large Raffel et al. (2020), zero-shot, few-shot, and Chain-of-Thought with GPT-4 OpenAI (2023), Llama 2 Touvron et al. (2023), etc. [^2] We also experiment with GPT-3.5, Flan-T5-XXL Chung et al. (2022), Alpaca-7B Taori et al. (2023) and Mistral-7B Jiang et al. (2023). We report the results in Appendix A.7.) through automated evaluation and human expert evaluations.
StorySparkQA can benefit different research directions in the children education domain, particularly in better understanding domain experts’ thinking process, and training models to generate story-based QA pairs infused with real-world knowledge, with the ultimate goal of broadening children’s knowledge scope beyond story narratives that parents expect. In addition, we believe our annotation framework has the potential to be generalized to domain-specific tasks analogous to the real world that require structured external knowledge Vrandečić and Krötzsch (2014); Lehmann et al. (2015), such that clinicians use structured guidelines and knowledge for diagnosing ElSayed et al. (2023); American Diabetes Association (2011).
2 Related Work
2.1 Children Education and Real-World Knowledge Resources
Existing datasets in the education domain (e.g., StoryQA Zhao et al. (2023), FairytaleQA Xu et al. (2022), and EduQG Hadifar et al. (2023)) mostly comprise QA-pairs grounded in the story, lacking real-world knowledge beyond the story. We present key properties of related children education datasets in Table 6. On the other hand, general-purpose datasets like CommonsenseQA Talmor et al. (2018) and SciQA Auer et al. (2023) integrate crowd-sourced commonsense with narratives, but lack educational appropriateness aligned with children’s knowledge level.
Many popular real-world knowledge resources, such as ATOMIC Sap et al. (2019) and Wikidata Vrandečić and Krötzsch (2014), are too complicated for children’s knowledge level. A more appropriate option is ConceptNet Speer et al. (2017), a very large-scale knowledge graph for real-world concepts and relations stored in triples: ($concept_{1}$, $relation$, $concept_{2}$). The simplicity of triple representations makes ConceptNet suitable for children education, as demonstrated in prior literature Xu et al. (2020), thus, our work also leverages ConceptNet to support experts’ annotation process.
2.2 QA Pair Annotation Frameworks
Some existing annotation frameworks, such as Potato Pei et al. (2022) and Piaf Keraron et al. (2020), mostly focus on facilitating extractive QA pairs grounded in the text, that is, providing source texts and allowing annotators to highlight a span of text as an answer to a question. Some others, like the annotation toolkit for StoryQA Zhao et al. (2023), support free-form input, allowing annotators to type in answers in their own words through the data collection user interface. In either type, existing annotation frameworks can’t support story-based external knowledge collection and story data annotation effectively, in which annotators are required to recall comprehensive and systematical real-world knowledge for a given story text. Our study bridges this gap by proposing an external knowledge-empowered annotation framework.
2.3 QA Pair Generation (QAG)
Fine-tuning traditional pre-trained language models like BERT Devlin et al. (2019) on QAG datasets for end-to-end generation was a prevalent approach, but such methodology heavily depends on the training data quality and lacks control of generated content, which is inappropriate for the children education domain. Existing works also attempted to design multi-step generation pipelines, which offer better control of the generated content Yao et al. (2021); Wan et al. (2024).
Recent advancement in large language models (LLMs), such as GPT-3.5, GPT-4 OpenAI (2023), and Llama 3 Dubey et al. (2024), supports free-form natural language input and output without the need for tuning model parameters. Many prompting strategies were developed to further enhance models’ task-solving and domain-adaptation capabilities, including few-shot in-context learning (i.e., add a few examples in input) Brown et al. (2020); Yao et al. (2024), Chain-of-Thought (i.e., ask models to think “step-by-step”) Wei et al. (2022), etc. The performance of these prompting methods on the general QAG task has also been evaluated Ling and Afzaal (2024); Lu et al. (2023). However, to what extent these prompting and modeling strategies are effective in the QAG task for knowledge beyond the story content, and whether a compact language model fine-tuned on domain-specific datasets performs better or worse than generic LLMs in the context of children education remains underexplored. For example, recent work demonstrates the unreliability of generic LLMs for the classification of mental health issues and the superiority of fine-tuning domain-specific language models with high-quality datasets Xu et al. (2024). Our work attempts to step forward through comprehensive evaluation in Section 5.
3 Expert Annotation Framework
To better understand experts’ mental procedures during annotating QA pairs enriched with external real-world knowledge, we proposed a three-step QA pair annotation framework with interactive user interfaces (UI). Particularly, considering the challenges facing annotators in recalling the comprehensive and systematical external knowledge for a given story text Berry et al. (2016), our framework incorporates ConceptNet, a large-scale real-world Knowledge Graph, to support experts’ large-coverage knowledge collection. The workflow of our annotation framework is shown in Figure 2.
Figure 2: Workflow of the experts’ annotation process. Experts need to select a concept first, then match it with the most suitable knowledge, and finally create a QA pair based on the selected knowledge.
Figure 3: The user interface to facilitate our annotation task. The words highlighted in grey are candidate concepts. The blue block shows the Wiktionary explanation, and the yellow block lists our recommended triples.
Step 1. Concept Selection
In this step, experts identify an educationally suitable concept from the story text. We develop a collection of heuristics to filter candidate concepts that are tier 1 or tier 2  [^3] Tier 1 words are common and basic words. Tier 2 contains high-frequency words of various domains Beck et al. (2013). vocabulary and a concrete noun, verb, or adjective. First, we leverage the spaCy Honnibal and Montani (2017) English model to filter auxiliary words and punctuation 444tagged by ‘auxiliary’, ‘adposition’, ‘determiner’, ‘particle’, ‘punctuation’, ‘symbol’, and ‘other’ from the original story text. Then, we use AllenNLP’s Gardner et al. (2017) semantic role labeling tool to tag the latent structure of each sentence in the story context. This process identifies and retains key elements represented by semantic roles, including agents, goals, and results, which are subsequently treated as potential candidate concepts. We design the UI for step 1 to display one story section and allow experts to select highlighted candidate concepts in grey, as shown in Figure 5.
Step 2. Knowledge Matching
This step allows experts to select real-world knowledge based on the concept selected previously. Inspired by  Xu et al. (2020)’s work of combining and filtering knowledge from Wiktionary  [^5] https://www.wiktionary.org/ and ConceptNet Speer et al. (2017) for commonsense question answering, we implement a knowledge matching module that can retrieve and rank external knowledge associated with each concept selected by the experts.
Specifically, once experts select a candidate concept, our knowledge matching module 1) retrieves a list of real-world knowledge triples, with the format of (source concept, relation, target concept) from ConceptNet; 2) filters out weak relations in ConceptNet (complete relation list in Appendix A.3), and 3) rank knowledge triples by concatenating concepts and relationships, and calculating the average similarity between every other triple with the Term Frequency-Inverse Document Frequency (TF-IDF) Ramos et al. (2003).
We rank all retrieved triples with $1-\overline{s}+w$, where $\overline{s}$ denotes the similarity score and $w$ denotes the weight of a triple provided by ConceptNet, reflecting the combined influence and credibility of the triple by summing up the weights coming from all the sources that support it. The top six ranked triples are shown to annotators to balance between providing a sufficient selection and avoiding excessive distractions during the annotation task. We also retrieve the explanation for expert-selected concepts from Wiktionary to better facilitate experts’ annotations. The UI for this step is designed to display a list of recommended real-world knowledge triples along with the Wiktionary explanation to facilitate experts’ selection (Figure 6).
Step 3. QA pair Annotation
This step enables annotators to create a QA pair based on the real-world knowledge triple they selected in step 2, and the corresponding UI is shown in Figure 3. In this step, experts are instructed to incorporate one concept in the question or answer and include the relation from the triple in the resulting QA pair.
4 StorySparkQA
{booktabs}
colspec=l|cccc|cccc|cccc, cells=valign=m, width=row1=font=, row3=font=, cell11=r=3c,m, cell12,6,10=c=4c, cell22,6,10=c=4c, cell4-102-13=font=, StorySparkQA Train Validation Test
$232$ books with $4,300$ QA pairs $23$ books with $769$ QA pairs $23$ books with $799$ QA pairs
Mean SD Min Max Mean SD Min Max Mean SD Min Max
# sections / story $14.4$ $8.8$ $2$ $60$ $16.5$ $10.0$ $4$ $43$ $15.8$ $10.8$ $2$ $55$
# tokens per story $2160.9$ $1375.9$ $228$ $7577$ $2441.8$ $1696.9$ $425$ $5865$ $2313.4$ $1369.6$ $332$ $6330$
# tokens / section $149.6$ $64.8$ $12$ $447$ $147.8$ $56.7$ $33$ $298$ $145.8$ $58.6$ $24$ $290$
# questions / story $18.5$ $14.5$ $2$ $126$ $33.4$ $22.1$ $4$ $115$ $34.7$ $21.1$ $8$ $90$
# questions / section $1.3$ $0.6$ $1$ $9$ $2.1$ $0.3$ $2$ $3$ $2.1$ $0.3$ $2$ $3$
# tokens / question $5.2$ $2.0$ $3$ $19$ $5.9$ $1.6$ $3$ $13$ $6.0$ $1.7$ $3$ $13$
# tokens / answer $5.4$ $3.7$ $1$ $20$ $3.8$ $2.3$ $1$ $12$ $3.8$ $2.3$ $1$ $12$
Table 1: Core statistics of our StorySparkQA dataset, which has $278$ books and $5,868$ QA pairs.
StorySparkQA aims to facilitate parents’ interactive story reading with appropriate real-world knowledge: practical, factual, everyday information that helps preschoolers understand the world around them. Our dataset consists of $5,868$ QA pairs annotated by children education experts leveraging our designed annotation framework. We present the core statistics of StorySparkQA in Table 4 and show one example in Figure 1.
4.1 Source Narrative
Among the existing story-based datasets for children education, FairytaleQA Xu et al. (2022) comprises $278$ classic fairytale stories of various origins, and all the stories have been evaluated as suitable for 8th-grade children and younger. The original stories were parsed by education experts into shorter sections of around $150$ words, which leads the FairytaleQA dataset to a unique and high-quality text corpus for children’s reading comprehension. As a result, we take the story sections from FairytaleQA as the source text for our StorySparkQA dataset.
4.2 Annotation Process
Following our annotation framework, we recruit 11 children education experts for the annotation task. The experts all have a minimum of 3 years of practical experience (e.g., kindergarten teachers) in learning science and possess relevant educational backgrounds. For each story section, experts are asked to first identify a concept from the story. Educational experts should select concepts that are most beneficial for children’s education from story text. The experts then proceed to select a real-world knowledge triple associated with the selected concept and create a QA pair based on the selected triple. In this process, experts are asked to take into account children’s cognitive and knowledge levels and write QA pairs that are most appropriate for 3- to 6-year-olds. We collect experts’ mental procedures by recording their selected concepts, real-world knowledge triples, and created QA pairs during the annotation process.
4.2.1 Cross-Validation
To ensure the quality and consistency of annotated QA pairs among annotators, as well as to evaluate agreement in selecting triples and creating QA pairs between annotators, we designed additional cross-validation procedures with corresponding UIs. We randomly selected $50$ QA pairs in both the test and validation split ($100$ QA pairs in total) and two annotators were asked to cross-validate each other’s annotation (denoted by $annotator_{A}$ and $annotator_{B}$, accordingly):
  1. 1.
    Shown in Figure 7, $annotator_{A}$ is provided with the story section and the concept selected by $annotator_{B}$. For each selected concept, $annotator_{A}$ is asked to rank the top 3 triples from the same recommended triple list given to $annotator_{B}$, verifying the triple selection agreement between annotators (Figure 8).
  2. 2.
    In the next step, $annotator_{A}$ is asked to create an QA pair based on the word and triple selected by $annotator_{B}$, evaluating the similarity of QA pairs between annotators given the identical triple (Figure 9).
  3. 3.
    After submitting the QA pair in Step $2$, $annotator_{A}$ is provided with the question created by $annotator_{B}$ based on the same triple, and $annotator_{A}$ is asked to write an answer to the question to cross-validate the question-answering agreement (Figure 10).
Of the $100$ randomly selected sections in the validation and test splits, $86\%$ of the triples that appear in the top-3 list are selected by both annotators, and $56\%$ of the triples are ranked top by the validator, indicating a very high consistency between experts for triple selection. In addition, we evaluate the similarity of the concatenated QA pairs created by each of the annotators based on the same triple with Rouge-L Lin (2004) and SBERT Reimers and Gurevych (2019) scores. The Rouge-L F1 score of QA pair creation between annotators is 0.53, and the SBERT score is 0.80, showing a shared tendency among experts when selecting real-world knowledge and creating a QA pair that is both beneficial and suitable for children education.
Figure 4: Distribution of real-world knowledge relations annotated by experts in the StorySparkQA dataset
4.3 Statistics and Analysis of StorySparkQA
Statistics of StorySparkQA
Figure 4 demonstrates the distribution of real-world knowledge relations in StorySparkQA, and Table 4 illustrates detailed statistics of the dataset. On average, each section is annotated with approximately 1.4 QA pairs. In StorySparkQA, the top 3 real-world knowledge relations selected by experts are “is a”, “has subevent” and “is the antonym of”, respectively constituting $35.5$%, $16.2$% and $15.2$% of all real-world knowledge relations. The distribution of question types in StorySparkQA is shown in Table A.4 in Appendix A.4. In StorySparkQA, questions starting with “what” are the most common type of question, constituting $86.0$%. Questions starting with “why” and “how” constitute about $7.2$% and $2.4$%, respectively.
Analysis of StorySparkQA’s Alignment with Real-World Needs
According to experts’ annotations, real-world knowledge relation “is a” and questions start with “what” have a much higher proportion than the others. Moreover, of all experts’ annotations, the selected concept words are mostly nouns (65.06%) and adjectives (22.26%), which are easy for children to comprehend. Considering the cognitive development of children, especially those aged 3 to 6, they are typically in the stage of language development and exploration, full of curiosity about the world Chouinard et al. (2007); Jirout and Klahr (2012). It is therefore natural for them to ask questions as a way to satisfy this curiosity. Consequently, parents are more inclined to use “what” questions with simple vocabulary to inspire children’s thinking and encourage them to actively acquire knowledge Taylor et al. (1994); Yu et al. (2019). Consistent with the actual habits of parents, experts’ annotated questions have a high consensus that “what” questions are more in line with children’s learning and cognitive characteristics.
Comparison with Existing Datasets
Compared with existing QA datasets for children education Xu et al. (2022); Zhao et al. (2023), StorySparkQA is unique mainly in the annotation process and data composition. Going beyond direct QA pair annotation utilized by most QA datasets (e.g., FairytaleQA and StoryQA), StorySparkQA’s annotation process provides step-by-step support with structured real-world knowledge to ease the effort that the experts need to craft children-appropriate and knowledgeable QA pairs. In addition, StorySparkQA includes not only QA pairs, but also corresponding story texts and expert-selected real-world knowledge triples, which reflects experts’ mental procedures when creating these children-appropriated, knowledgeable QA pairs and leads to a more comprehensive dataset for children’s knowledge expansion. Catering to parents’ practical needs for knowledge expansion in interactive story reading, StorySparkQA complements existing QA datasets, which often focus on narrative comprehension and commonsense question-answering. The key properties of StorySparkQA and related children education datasets are illustrated in Table 6 in Appendix A.2.
5 Benchmark Experiment
We benchmark the quality and usability of our StorySparkQA on the QAG task, which is required to meet the needs of parents to guide children to learn some real-work knowledge during practical interactive story reading, as well as existing work of developing AI-assisted storytelling and reading systems Yao et al. (2021); Dietz et al. (2021); Zhang et al. (2022). We conduct an automated evaluation, reported in Section 5.1 to measure the semantic similarity of generated QA pairs with experts-annotated QA pairs, benchmarked with a T5-Large model fine-tuned on StorySparkQA and a set of robust LLMs. Considering the limitation of automated evaluation in evaluating the educational appropriateness of generated QA pairs, we further conduct a human evaluation, reported in Section 5.2, with children education experts.
5.1 Automated Evaluation
We now elaborate on the settings and results of our QAG experiments with various language models, through which to demonstrate the usability of StorySparkQA.
5.1.1 Experiment Settings
The QAG task involves taking a story section as input and generating the QA pairs. To exploit LLMs’ comprehensive generation ability, we design two variations to simulate experts’ annotation process:
  1. 1.
    QA pair generation: Generate the QA pair.
  2. 2.
    QA pair and triple generation: Generate the associated real-world knowledge triple alongside the QA pair.
The automatic evaluation comprises six popular LLMs: GPT-3.5, GPT-4 OpenAI (2023), FLAN-T5-XXL Chung et al. (2022), Alpaca-7B Taori et al. (2023), Mistral-7B Jiang et al. (2023) and Llama 2(7B) Touvron et al. (2023). We carefully design the prompt inputs (Appendix A.8) with clear and informative instructions, including 13 relation types (Appendix A.3) in ConceptNet. The goal is to leverage LLMs to generate diverse triples similar to those created by human education experts.
For each LLM involved in this experiment (GPT-3.5, GPT-4, FLAN-T5-XXL, Alpaca, Mistral, and Llama 2), we employ zero-shot, few-shot in-context learning (ICL) Brown et al. (2020) approaches to thoroughly examine the QAG performance of these models with different prompting strategies. For GPT-3.5 and GPT-4, we also use Chain-of-Thought (CoT) Wei et al. (2022) to further explore their QAG capabilities. Randomly sampled examples from the validation split are used as demonstrations for the few-shot ICL approaches. We also fine-tune a T5-Large model to examine how a much smaller domain-specific model, supported by expert-annotated triples as additional input, performs compared to generic LLMs. The experiment settings and hyper-parameters are reported in Appendix A.6.
We utilize Rouge-L Lin (2004) to evaluate the quality of concatenated QA pairs between the generated ones and two expert-annotated ground truths of each data, and report the averaged score across all data in the test split. For the setting that generates triples along with QA pairs, we evaluate the generated triples and QA pairs separately. The Rouge-L F1 score is chosen because it captures the sequential matching of texts, reflecting whether the generated QA pairs are textually similar to experts’ annotation. However, we are aware that Rouge-L is not without limitations, such as a lack of semantic understanding. To address this, we incorporate SBERT using Sentence Transformer Reimers and Gurevych (2019) for automated evaluation and present the full results in Appendix A.7. We also conduct a human evaluation (see in Section 5.2) to further assess the quality of generated QA pairs. We perform experiments with GPT-3.5 and GPT-4 three times for each setting to calculate a robust and reliable average score.
5.1.2 Results and Analysis
{booktabs}
colspec=cl|cc, row1=font=, width=cells=m, Model Prompting
Strategy QAG
w/o Triples QAG
w/ Triples
T5-Large
Fine-Tuned
(0.77B) - 0.332 0.279
[…]
(7B) zero-shot 0.124 0.266
few-shot 0.251 0.239
[…]
(7B) zero-shot 0.229 0.209
few-shot 0.267 0.257
[…]
(7B) zero-shot 0.213 0.177
1-shot 0.192 0.206
5-shot 0.241 0.269
[…]
1-shot 0.239 0.252
5-shot 0.262 0.264
CoT - 0.259
[…]
1-shot 0.272 0.251
5-shot 0.287 0.248
CoT - 0.262
Table 2: QAG performance of LLMs with different prompting strategies and the fine-tuned T5-Large model. Bolded numbers are the best scores within each setting.
In table 5.1.2, we show the zero-shot, few-shot ICL, and CoT performances on all models in both settings of the QAG task.
Generally, zero-shot QAG performance on these models falls short of the few-shot ICL QAG performance. Particularly, Alpaca’s few-shot performance when generating QA pairs without triples falls behind the zero-shot setting. This is due to the repetitive and less diverse outputs generated by the model in the zero-shot setting (e.g., ‘What is the relation between A and B? A is the antonym of B.’ ). Such QA pairs are repetitive across the results, leading to a higher Rouge-L score as the words in it have a higher chance to match with the experts’ annotations. Particularly, the relation ‘is the antonym of’ occupies a significant proportion in the concatenation of QA pairs.
Remarkably, models using 5-shot demonstrations outperform those using 1-shot demonstrations. It is notable that GPT-3.5 achieves better performance in the few-shot setting compared to GPT-4. We believe this is caused by GPT-4’s exceptional Natural Language Generation versatility, potentially resulting in longer and more grammatically complicated QA pairs that are unsuitable for 3- to 6-year-olds due to advanced vocabulary and knowledge comprehension challenges Ouellette (2006); Perfetti and Stafura (2014). Meanwhile, models employing the Chain-of-Thought prompting method do not imply an obvious improvement compared to the few-shot ICL QAG performance.
For the setting of generating triples along with QA pairs (w/ triples), the results do not indicate an improvement in QAG through the step of generating real-world knowledge triples. We attribute this to the potential complexity of the task that asks LLMs to generate real-world knowledge triples and corresponding QA pairs simultaneously.
It is worth noting that T5-Large fine-tuned on our StorySparkQA has a relatively better performance in generating QA pairs enriched with real-world knowledge for children than conversational LLMs like GPT-3.5 and GPT-4 by Rouge-L. Additional analysis of the generated external knowledge types among experts’ annotations, fine-tuned T5-Large, and GPT-4 is demonstrated in Appendix A.5.
5.2 Human Evaluation
{booktabs}
colspec=llccccc, row1=font=, width= Dimension Model Mean SD t df p-value
[…]
T5-Large Fine-Tuned 4.842 0.585 1.259 279 0.209
GPT-4 4.871 0.514 0.646 279 0.519
[…]
T5-Large Fine-Tuned 4.329 1.111 5.487 279 <0.01
GPT-4 4.379 0.869 5.123 279 <0.01
[…]
T5-Large Fine-Tuned 4.639 0.972 5.487 279 0.729
GPT-4 4.529 0.974 2.240 279 0.026
[…]
Appropriateness** T5-Large Fine-Tuned 4.325 0.972 2.937 279 <0.01
GPT-4 4.318 2.974 3.113 279 <0.01
Note: * denotes p-value <0.05, and ** denotes p-value <0.01. Both of the denotations indicate results are statistically significant.
Table 3: The paired sample t-test result of children education experts in comparison of GPT-4 and T5-Large fine-tuned on StorySparkQA in the QAG task. Bolded numbers are the best scores within each dimension excluding human experts’ annotations.
To further compensate for the limitation of Rouge-L and SBERT as well as to thoroughly assess the quality and usability of LLM-generated QA pairs, particularly in terms of educational appropriateness, we conducted a human study with four education experts to compare expert-annotated QA pairs and those generated by fine-tuned T5-Large and GPT-4 with 5-shot ICL, the best-performing ones in automated evaluation.
We randomly select ten story books from the test split of StorySparkQA, and sample seven sections per book. For each section, three QA pairs are created based on the story narrative (experts’ annotations, and QA pairs generated by GPT-4 and fine-tuned T5-Large), summing up $210$ QA pairs for the human evaluation. QA pairs are randomized for each section, and the sources are omitted to the human subjects for a fair evaluation.
Four experts evaluate each QA pair on the following four dimensions with a 5-point Likert scale:
  1. 1.
    Grammar Correctness: The QA pair uses comprehensible English Grammar;
  2. 2.
    Answer Relevancy: The answer is correct and corresponds to a question;
  3. 3.
    Contextual Consistency: The QA pair originates from the story and goes beyond the story’s immediate context;
  4. 4.
    Children’s Educational Appropriateness: The QA pair is appropriate for children’s reading experience during interactive story reading.
5.2.1 Results and Analysis
Table 5.2 illustrates the average scores of each dimension and paired sample t-test results. We observe that expert-created QA pairs outperform those generated by models in all four dimensions. The paired sample t-test results show that experts’ annotations are significantly different in three out of four dimensions compared with models’ generation. These justify StorySparkQA’s utility in catering to parents’ real-world needs in interactive story reading.
In terms of Grammar Correctness and Answer Relevancy, GPT-4 achieves better performance than the fine-tuned T5-Large. We believe it to be reasonable because LLMs such as GPT-4 are trained on vast amounts of corpora, enabling them to generate QA pairs with greater consistency in word usage. Therefore, compared with T5-Large, GPT-4 produces answers that connect more closely with the questions, resulting in greater coherence and accuracy between the questions and the answers.
In terms of Contextual Consistency, the fine-tuned T5-Large significantly outperformed GPT-4, behind experts’ annotations. A similar result could be found in Children’s Educational Appropriateness, wherein the T5-Large model fine-tuned on StorySparkQA also exhibits better performance.
These results suggest that fine-tuned with experts’ annotations, the T5-Large model can generate QA pairs that 1) contain external structured knowledge connected to the story narrative, and 2) are appropriate for young children to learn during the interactive story reading activities.
5.3 Discussion
Comparing the best-performing SoTA LLMs in the QAG pipeline with the corresponding fine-tuned T5-Large, we can observe that the T5-Large can reliably generate QA pairs aligned more with experts’ annotations in terms of Rouge-L score according to system evaluation, regardless of whether generating QA pairs along real-world knowledge triples. Drawing from the results of our human evaluation, the fine-tuned T5-Large exhibits better capabilities in generating QA pairs that suit parents’ real-world educational expectations of interactive story reading: originating from the story and embodying educational-appropriate real-world knowledge. Worth mentioning that T5-Large only consists of $770$ million parameters, whereas Alpaca-7B, Mistral-7B, and Llama 2 in our experiments consist of $7$ billion parameters ($10$ times larger).
This observation justifies StorySparkQA’s utility in training a task-specific model that caters to parents’ real-world story reading needs on the one hand, and demonstrates the usefulness of combining structured real-world knowledge and free-form narratives in domain-specific tasks such as interactive story reading.
6 Conclusion and Future Work
In summary, we propose StorySparkQA, an expert-annotated, external-knowledge-enriched QA dataset for children education, by leveraging a novel annotation framework to facilitate scalable expert annotations through structured external knowledge. We demonstrate the effectiveness of StorySparkQA through an automated evaluation on various LLMs of generating QA pairs catering to parents’ needs and a human evaluation with children education experts.
One possible future work is refining the QAG pipeline structures and exploiting LLMs to generate QA pairs that align more closely with parents’ practical needs. Another future direction involves using StorySparkQA and language models to develop a human-AI collaborative education system (e.g., an interactive story reading system), aiding parents and educators to formulate personalized questions during story readings, while addressing their language, knowledge, or time constraints. Also, fine-tuning LLMs (e.g., Llama 3) may lead to better performance on the QAG task, which offers a future direction to refine models’ capabilities in real-world tasks like children’s education.
7 Acknowledgement
We thank the Lab of Artificial Intelligence for Education, East China Normal University, Shanghai Institute of Artificial Intelligence for Education, East China Normal University, and Shanghai Science and Technology Innovation Action Plan project (No. 21511104500) for funding this research.
8 Limitations
This work primarily focuses on constructing an expert-annotated, large-scale QA dataset consisting of story-based QA pairs associated with real-world knowledge beyond the story narrative, however, this work is not without limitations. We could further explore LLMs’ QAG capabilities with different models (e.g., GPT-4 and Llama 3), and domain-specific prompting methodologies, such as ICL with more demonstrations and RAG approaches Edge et al. (2024) with multi-step generation pipelines. In addition, the size and scope limitations of expert annotations in our dataset may not be sufficient for developing NLP technologies that can be generalizable to similar scenarios. We call for future research to explore methods for scaling the data annotation process in real-world settings and to investigate strategies for efficiently optimizing or evaluating NLP technologies in low-resource scenarios where expert resources are scarce.
Appendix A Appendix
A.1 Sample Data of StorySparkQA
In Table A.1 and  A.1, we present the sample data from StorySparkQA, which include expert-selected concepts, real-world knowledge triples, and created QA pairs.
{booktabs}
width=colspec=Q[l,co=1,m], Story Section:
At the time when the Tang dynasty reigned over the Middle Kingdom, there were master swordsmen of various kinds.
Those who came first were the saints of the sword. They were able to take different shapes at will, and their swords were like strokes of lightning.
They wore a hidden dagger at their side and carried a leather bag at their belt.
By magic means they were able to turn human heads into flowing water.
Expert annotated QA pairs
Triple: (dagger, is a, short sword)
Question: What is a short sword called?
Answer: A dagger.
Triple: (bag, is used for, carrying things)
Question: What is a bag used for?
Answer: A bag is used for carrying things.
Table 4: Example 1 of expert annotated data point in StorySparkQA.
{booktabs}
width=colspec=Q[l,co=1,m], Story Section:
On hearing this the king walked to the window and stood for a few minutes with his back to the room, where the company of young men remained silent. Then he came back, his face white and stern.
’I tell you,’ he said, ’and it is the solemn truth, that I would rather you had told me that the prince was dead, though he is my only son, than know that he would suffer such an injury without attempting to avenge it
Expert annotated QA pairs
Triple: (white, is a, color)
Question: What color is snow?
Answer: A White.
Triple: (injury, is at location of, hospital)
Question: Where do you go if you get very hurt?
Answer: You go to hospital if you get very hurt.
Table 5: Example 2 of expert annotated data point in StorySparkQA.
A.2 Properties of Educational QA datasets
The key properties of educational QA datasets, including their number and type of sourcebooks, QA pairs, whether they contain external knowledge, annotators, annotation process, and data composition, are presented in Table 6.
Dataset # Books # QA Pairs External Knowledge Annotator Document Source Annotation Process Data Composition
StoryQA 148 38,703 Yes Crowd-Sourced Story books Direct Annotation 1. Story Section 2. QA Pairs
FairytaleQA 278 10,580 No Expert Story books Direct Annotation 1. Story Section 2. QA Pairs
EduQG 13 5,018 No Expert Text books Direct Annotation 1. Source Documents 2. Questions and Answer Options
StorySparkQA 278 5,868 Yes Expert Story books 3-step Guided Annotation 1. Story Section 2. Real-world Knowledge Triples 3. QA Pairs
Table 6: Properties of existing datasets focusing on children education compared with our StorySparkQA.
A.3 ConceptNet Relations
We follow Xu et al. (2020)’s work to filter out weak relations in ConceptNet, and our ranking algorithm uses the following 13 relations in our annotation framework as well as GPT prompts: causes, desires, has context of, has property, has subevent, is a, is at location of, is capable of, is created by, is made of, is part of, is the antonym of, is used for.
A.4 Distribution of Question Type
The distribution of question type in StorySparkQA is shown in Table A.4.
{booktabs}
colspec=ccccr, row1=font=, cells=m, Interrogative Train
Split Val
Split Test
[…]
Percentage (%)
what 3779 628 641 86.01
why 227 93 105 7.24
who 76 10 14 1.70
where 41 3 7 0.87
when 20 12 8 0.68
how 112 13 15 2.39
other 42 10 9 1.04
Table 7: Distribution of question types in StorySparkQA.
A.5 Analysis of the Type of Generated External Knowledge
We calculated the real-world knowledge type distribution of experts’ annotation and triples generated by T5-Large and GPT-4 with 5-shot ICL (the best-performing ones in automated evaluation). The results are shown in Table 8.
According to the result, the fine-tuned T5 model can generate real-world knowledge triples more aligned with experts’ annotation, surpassing GPT-4, further proving that domain-specific fine-tuning provides the model with targeted knowledge and expertise that LLMs may lack.
Relations Experts’ Annotation T5-Large Fine-Tuned GPT-4
is a 35.45% 47.67% 20.64%
has subevent 16.21% 16.44% 4.66%
is the antonym of 15.20% 7.95% 1.64%
is used for 8.78% 9.32% 35.25%
is at location of 7.53% 7.40% 2.19%
is capable of 5.18% 4.11% 9.77%
other 11.65% 6.85% 25.84%
Table 8: Comparison of real-world knowledge relation types across experts’ annotations, fine-tuned T5-Large, and GPT-4.
A.6 Hyper-Parameters and Experiment Settings
We conducted our experiments on Google Colab with A100. Following common practice when fine-tuning the T5-Large model, we use the learning rate of 1e-4 and train our model on 3 epochs.
A.7 Complete QAG Pipeline Results
We demonstrate the complete performance of LLMs in our QAG pipeline using both zero-shot and few-shot ICL approaches in Table A.7.
{booktabs}
[…]
w/ Triples
Rouge-L SBERT Rouge-L SBERT
T5-Large
Fine-Tuned (0.77B) - 0.332 0.289 0.279 0.263
[…]
(7B) zero-shot 0.124 0.186 0.266 0.207
1-shot 0.251 0.182 0.239 0.186
[…]
(7B) zero-shot 0.229 0.237 0.209 0.229
1-shot 0.227 0.237 0.231 0.241
5-shot 0.267 0.241 0.257 0.251
[…]
(7B) zero-shot 0.213 0.234 0.177 0.225
1-shot 0.192 0.217 0.206 0.237
5-shot 0.241 0.240 0.269 0.253
Flan-T5-XXL 1-shot 0.264 0.246 0.194 0.209
[…]
1-shot 0.239 0.262 0.252 0.271
5-shot 0.262 0.279 0.264 0.266
CoT - - 0.259 0.280
[…]
1-shot 0.272 0.279 0.251 0.292
5-shot 0.287 0.311 0.248 0.283
CoT - - 0.262 0.292
Table 9: Rouge-L and SentenceBERT scores of LLMs in the QAG task. Bolded numbers are global best performance within each setting on each metric.
A.8 LLM Prompts
To utilize LLMs’ strong reasoning and generation capability as well as control GPT-generated questions as much as possible to meet the needs of parents, we carefully design our prompts.
For the QAG pipeline, there are two variations based on the system: (1) Directly generate a QA pair based on a provided story section. (2) From a story section, generate a real-world knowledge triple and a QA pair simultaneously.
Table A.8A.8 list our prompts for GPT in the two abovementioned approaches.
{booktabs}
width=colspec=Q[l,co=1,m], cell11=c, Prompt for LLMs in the QAG Pipeline
(Generate QA Pairs Only)
I need you to help generate a question and answer pair for young children aged three to six. I will provide you with a short section of a story delimited by triple quotes. Please follow these steps:
1. For each sentence, identify one key word that meets the following criteria: it is relatively complex, it is considered tier 1 or tier 2 vocabulary, and it is a concrete noun, verb, or adjective.
2. After this, you need to completely forget about the story that I gave you, remembering only the words you identified.
3. Based on each selected word, generate a question and answer pair that either the question or the answer contains that word. For example, if your identified word is ’apple’, your question could be: where do apples grow? what do apples taste like? what color are apples? These questions should go beyond the context of the stories.
Each question should have one single correct answer that would be the same regardless of the children’s experiences. The questions should be focused on real-world, fact-based knowledge and beneficial to educate children during story reading.
The real-world, fact-based knowledge should be based on the selected word and is in the form of a triple such as A relation B, where A and B are two concepts and the selected word can be either A or B. You should use one of the following relations for the real-world knowledge:
    1) causes, 2) desires, 3) has context of, 4) has property, 5) has subevent, 6) is a, 7) is at location of,
    8) is capable of, 9) is created by, 10 ) is made of, 11) is part of, 12) is the antonym of, 13) is used for
4. After this, select one question-answer pair that you think best meets my criteria. Please note that the question should be answerable without reading the story. The answer should only be a concrete noun, verb, or adjective.
Return the selected question-answer pair in the following format:
question: …
answer: …
⟨story⟩:
{story1 for few-shot}
⟨response⟩:
{response1 for few-shot}
⟨story⟩:
{story for the current data}
⟨response⟩:
Table 10: Prompt for LLMs in the QAG task with generating QA pairs directly from the story.
{booktabs}
width=colspec=Q[l,co=1,m], cell11=c, Prompt for LLMs in the QAG Pipeline
(Generate Triples and QA Pairs)
I need you to help generate a question and answer pair for young children aged three to six. I will provide you with a short section of a story delimited by triple quotes. Please follow these steps:
1. For each sentence, identify one key word that meets the following criteria: it is relatively complex, it is considered tier 1 or tier 2 vocabulary, and it is a concrete noun, verb, or adjective.
2. After this, you need to completely forget about the story that I gave you, remembering only the words you identified.
3. Based on each selected word, generate one real-world relation based on the selected word. This real-world relation should go beyond the context of the stories. For example, if your identified word is ’apple’, your real-world relation could be: apple grows on trees; apples are red. The real-world, fact-based knowledge should be based on the selected word and is in the form of a triple such as ’A relation B’, where A and B are two concepts and the selected word can be either A or B. You should use one of the following relations for the real-world knowledge:
    1) causes, 2) desires, 3) has context of, 4) has property, 5) has subevent, 6) is a, 7) is at location of,
    8) is capable of, 9) is created by, 10 ) is made of, 11) is part of, 12) is the antonym of, 13) is used for
4. After this, generate a question and answer pair based on the real-world, fact-based knowledge you generated. Either the question or the answer should contain that identified word. Each question should have one single correct answer that would be the same regardless of the children’s experiences. The questions should be focused on real-world, fact-based knowledge and beneficial to educate children during story reading.
5. After this, select one question-answer pair that you think best meets my criteria. Please note that the question should be answerable without reading the story. The answer should only be a concrete noun, verb, or adjective.
Return the generated real-world knowledge triple and selected question-answer pair in the following format:
real-world knowledge triple: (A, relation, B)
question: …
answer: …
⟨story⟩:
{story1 for few-shot}
⟨response⟩:
{response1 for few-shot}
⟨story⟩:
{story for the current data}
⟨response⟩:
Table 11: Prompt for LLMs in the QAG task with generating real-world knowledge triple and QA pairs directly from the story.
A.9 User Interface for Annotation System
We implement an annotation system to facilitate QA pair annotation with associated external knowledge. Figure 56 and 3 show the annotation interface for human experts.
We also conduct cross-validation to assess the agreement among annotators. Figure 789 and 10 demonstrate user interfaces for each step to support the cross-validation process.
Figure 5: Annotation process 1: browse a displayed section, with candidate words highlighted in grey.
Figure 6: Annotation process 2: after selecting a word (highlighted in red), related explanation in Wiktionary and candidate real-world knowledge triples in ConceptNet will display.
Figure 7: Cross-validation process 1: browse a displayed section, with candidate words highlighted in grey.
Figure 8: Cross-validation process 2: select a word annotated by others and rank the candidate triples.
Figure 9: Cross-validation process 3: after ranking top3 triples, the triple selected originally by the other annotator is displayed, the validator should create a QA pair based on the original triple.
Figure 10: Cross-validation process 4: validator is asked to answer the question created by the other annotator using the triple originally selected by the other annotator.

### v3

@@
 Abstract
-Interactive story reading is a common parent-child activity, where parents expect to teach both language skills and real-world knowledge beyond the story.
-While increasing storytelling and reading systems have been developed for this activity, they often fail to infuse real-world knowledge into the conversation.
+Interactive story reading is common in early childhood education, where teachers expect to teach both language skills and real-world knowledge beyond the story.
+While many story reading systems have been developed for this activity, they often fail to infuse real-world knowledge into the conversation.
 This limitation can be attributed to the existing question-answering (QA) datasets used for children’s education, upon which the systems are built, failing to capture the nuances of how education experts think when conducting interactive story reading activities.
@@
 StorySparkQA: Expert-Annotated QA Pairs with Real-World Knowledge for Children’s Story-Based Learning
-Jiaju Chen East China Normal University                        Yuxuan Lu Northeastern University                        Shao Zhangthanks: Work done when visiting Northeastern University.
+Jiaju Chen* East China Normal University                        Yuxuan Lu Northeastern University                        Shao Zhangthanks: Work done when visiting Northeastern University.
 Shanghai Jiao Tong University
@@
 Yunyao Li Adobe                        Qianwen Wang University of Minnesota Twin Cities                        Dakuo Wang Northeastern University
-Yuling Sunthanks:  Corresponding Author: ylsun@cs.ecnu.edu.cn. East China Normal University
+Yuling Sunthanks:  Corresponding Author: ylsun@cs.ecnu.edu.cn. Lab of Artificial Intelligence for Education East China Normal University
 1 Introduction
@@
 Additional data examples of StorySparkQA are presented in Appendix A.1.
-Interactive story reading is a common parent-child activity, where parents often sit together with preschool children, read storybooks, and proactively engage in question-answering (QA) conversations with them Wright (1995); Isbell et al.
+Interactive story reading is common in early childhood education, where teachers often sit together with preschool children, read storybooks, and proactively engage in question-answering (QA) conversations with them Wright (1995); Isbell et al.
 (2004).
-Such guided conversations are typically grounded in but beyond the story narratives Kotaman (2013), with parents’ expectations of guiding children to learn real-world knowledge and improving their historical, cultural, and emotional awareness Sun et al.
+Such guided conversations are typically grounded in but beyond the story narratives Kotaman (2013), with teachers’ expectations of guiding children to learn real-world knowledge and improving their historical, cultural, and emotional awareness Sun et al.
 (2024).
@@
 (2021), etc.
-Despite the benefits of interactive story reading, parents often struggle to appropriately conduct such interactive story reading with children because of multi-facet difficulties Golinkoff et al.
+Despite the benefits of interactive story reading, teachers often struggle to appropriately conduct such interactive story reading with children because of multi-facet difficulties Golinkoff et al.
 (2019); Sun et al.
 (2024).
-Specifically, such interactive story reading needs parents to identify the knowledge of interest during storytelling, formulate the real-world knowledge piece they want to teach in mind ("what to ask"), then ask an engaging question ("how to ask") to children at the appropriate time ("when to ask").
-Yet, most parents lack the educational expertise necessary to guide such educational conversations Golinkoff et al.
+Specifically, such interactive story reading needs teachers to identify the knowledge of interest during storytelling, formulate the real-world knowledge piece they want to teach in mind ("what to ask"), then ask an engaging question ("how to ask") to children at the appropriate time ("when to ask").
+In home settings, most parents also lack the educational expertise necessary to guide such educational conversations Golinkoff et al.
 (2019); Sun et al.
 (2024).
-Also, parents in contemporary society often hardly maintain high concentration to accompany their children due to the need to deal with other work and family chores at the same time Zhang et al.
+Meanwhile, today’s parents often hardly maintain constant focus on their children due to the need to deal with other work and family chores at the same time Zhang et al.
 (2022); Sun et al.
@@
 (2023), MatheMyths Zhang et al.
-(2024)), have demonstrated effectiveness in children’s storytelling scenarios Dietz et al.
+(2024)), have demonstrated utility in children’s storytelling scenarios Dietz et al.
 (2021).
-These systems primarily take advantage of the verbal communication interface and advanced language models to support natural conversation with humans Mahmood et al.
+These systems primarily utilize the verbal communication interface and advanced language models to support natural conversation with humans Mahmood et al.
 (2023); Chan et al.
@@
 (2021).
-Nevertheless, existing AI-assisted storytelling and reading systems are not without limitations.
+Nevertheless, existing AI-assisted storytelling systems are not without limitations.
 Particularly, building on top of data resources with mostly extractive QA pairs (e.g., FairytaleQA Xu et al.
-(2022)) – where the answers can be found directly in the story narrative – these systems fall short in helping parents teach real-world knowledge beyond the story narrative Yao et al.
-(2021), which is one of the main expectations of parents Sun et al.
+(2022)) – where the answers can be found directly in the story narrative – these systems fall short in helping teaching real-world knowledge beyond the story narrative Yao et al.
+(2021); Zhao et al.
+(2022), which is one of the main expectations of parents and teachers Sun et al.
 (2024).
@@
 We report the results in Appendix A.7.) through automated evaluation and human expert evaluations.
-StorySparkQA can benefit different research directions in the children education domain, particularly in better understanding domain experts’ thinking process, and training models to generate story-based QA pairs infused with real-world knowledge, with the ultimate goal of broadening children’s knowledge scope beyond story narratives that parents expect.
+StorySparkQA can benefit different research directions in the children education, particularly in better understanding domain experts’ thinking process, and training models to generate story-based QA pairs infused with real-world knowledge, with the ultimate goal of broadening children’s knowledge scope beyond story narratives that parents and teachers expect.
 In addition, we believe our annotation framework has the potential to be generalized to domain-specific tasks analogous to the real world that require structured external knowledge Vrandečić and Krötzsch (2014); Lehmann et al.
@@
 Step 1. Concept Selection
-In this step, experts identify an educationally suitable concept from the story text.
+In this step (UI shown in Figure 5), experts identify an educationally suitable concept from the story text.
 We develop a collection of heuristics to filter candidate concepts that are tier 1 or tier 2  [^3] Tier 1 words are common and basic words.
@@
 This process identifies and retains key elements represented by semantic roles, including agents, goals, and results, which are subsequently treated as potential candidate concepts.
-We design the UI for step 1 to display one story section and allow experts to select highlighted candidate concepts in grey, as shown in Figure 5.
 Step 2. Knowledge Matching
-This step allows experts to select real-world knowledge based on the concept selected previously.
+This step (UI shown in Figure 6) allows experts to select real-world knowledge based on the concept selected previously.
 Inspired by  Xu et al.
@@
 We rank all retrieved triples with $1-\overline{s}+w$, where $\overline{s}$ denotes the similarity score and $w$ denotes the weight of a triple provided by ConceptNet, reflecting the combined influence and credibility of the triple by summing up the weights coming from all the sources that support it.
-The top six ranked triples are shown to annotators to balance between providing a sufficient selection and avoiding excessive distractions during the annotation task.
-We also retrieve the explanation for expert-selected concepts from Wiktionary to better facilitate experts’ annotations.
-The UI for this step is designed to display a list of recommended real-world knowledge triples along with the Wiktionary explanation to facilitate experts’ selection (Figure 6).
+The top six ranked triples are shown to annotators to balance between providing a sufficient selection and avoiding excessive distractions during annotation.
+We also retrieve the explanation for concepts from Wiktionary to better facilitate experts’ annotations.
 Step 3. QA pair Annotation
@@
 Table 1: Core statistics of our StorySparkQA dataset, which has $278$ books and $5,868$ QA pairs.
-StorySparkQA aims to facilitate parents’ interactive story reading with appropriate real-world knowledge: practical, factual, everyday information that helps preschoolers understand the world around them.
+StorySparkQA aims to facilitate teachers’ interactive story reading with appropriate real-world knowledge: practical, factual, everyday information that helps preschoolers understand the world around them.
 Our dataset consists of $5,868$ QA pairs annotated by children education experts leveraging our designed annotation framework.
@@
 The experts all have a minimum of 3 years of practical experience (e.g., kindergarten teachers) in learning science and possess relevant educational backgrounds.
-For each story section, experts are asked to first identify a concept from the story.
-Educational experts should select concepts that are most beneficial for children’s education from story text.
+For each story section, experts are asked to first identify a concept from the story by selecting concepts that are most beneficial for children’s education from story text.
 The experts then proceed to select a real-world knowledge triple associated with the selected concept and create a QA pair based on the selected triple.
-In this process, experts are asked to take into account children’s cognitive and knowledge levels and write QA pairs that are most appropriate for 3- to 6-year-olds.
+In this process, experts are asked to consider children’s cognitive and knowledge levels and write QA pairs that are most appropriate for 3- to 6-year-olds.
 We collect experts’ mental procedures by recording their selected concepts, real-world knowledge triples, and created QA pairs during the annotation process.
@@
 It is therefore natural for them to ask questions as a way to satisfy this curiosity.
-Consequently, parents are more inclined to use “what” questions with simple vocabulary to inspire children’s thinking and encourage them to actively acquire knowledge Taylor et al.
+Consequently, teachers are more inclined to use “what” questions with simple vocabulary to inspire children’s thinking and encourage them to actively acquire knowledge Taylor et al.
 (1994); Yu et al.
 (2019).
-Consistent with the actual habits of parents, experts’ annotated questions have a high consensus that “what” questions are more in line with children’s learning and cognitive characteristics.
+Consistent with the actual habits of teachers, experts’ annotated questions have a high consensus that “what” questions are more aligned with children’s cognitive characteristics.
 Comparison with Existing Datasets
@@
 In addition, StorySparkQA includes not only QA pairs, but also corresponding story texts and expert-selected real-world knowledge triples, which reflects experts’ mental procedures when creating these children-appropriated, knowledgeable QA pairs and leads to a more comprehensive dataset for children’s knowledge expansion.
-Catering to parents’ practical needs for knowledge expansion in interactive story reading, StorySparkQA complements existing QA datasets, which often focus on narrative comprehension and commonsense question-answering.
+Catering to teachers’ practical needs for knowledge expansion in interactive story reading, StorySparkQA complements existing QA datasets, which often focus on narrative comprehension and commonsense question-answering.
 The key properties of StorySparkQA and related children education datasets are illustrated in Table 6 in Appendix A.2.
 5 Benchmark Experiment
-We benchmark the quality and usability of our StorySparkQA on the QAG task, which is required to meet the needs of parents to guide children to learn some real-work knowledge during practical interactive story reading, as well as existing work of developing AI-assisted storytelling and reading systems Yao et al.
+We benchmark the quality and usability of our StorySparkQA on the QAG task, which is required to meet the needs of teachers to guide children to learn some real-work knowledge during practical interactive story reading, as well as existing work of developing AI-assisted storytelling and reading systems Yao et al.
 (2021); Dietz et al.
@@
 GPT-4 4.318 2.974 3.113 279 <0.01
-Note: * denotes p-value <0.05, and ** denotes p-value <0.01. Both of the denotations indicate results are statistically significant.
 Table 3: The paired sample t-test result of children education experts in comparison of GPT-4 and T5-Large fine-tuned on StorySparkQA in the QAG task.
 Bolded numbers are the best scores within each dimension excluding human experts’ annotations.
+* means p-value <0.05, and ** means p-value <0.01, both are statistically significant.
 To further compensate for the limitation of Rouge-L and SBERT as well as to thoroughly assess the quality and usability of LLM-generated QA pairs, particularly in terms of educational appropriateness, we conducted a human study with four education experts to compare expert-annotated QA pairs and those generated by fine-tuned T5-Large and GPT-4 with 5-shot ICL, the best-performing ones in automated evaluation.
@@
 The paired sample t-test results show that experts’ annotations are significantly different in three out of four dimensions compared with models’ generation.
-These justify StorySparkQA’s utility in catering to parents’ real-world needs in interactive story reading.
+These justify StorySparkQA’s utility in catering to teachers’ real-world needs in interactive story reading.
 In terms of Grammar Correctness and Answer Relevancy, GPT-4 achieves better performance than the fine-tuned T5-Large.
@@
 Comparing the best-performing SoTA LLMs in the QAG pipeline with the corresponding fine-tuned T5-Large, we can observe that the T5-Large can reliably generate QA pairs aligned more with experts’ annotations in terms of Rouge-L score according to system evaluation, regardless of whether generating QA pairs along real-world knowledge triples.
-Drawing from the results of our human evaluation, the fine-tuned T5-Large exhibits better capabilities in generating QA pairs that suit parents’ real-world educational expectations of interactive story reading: originating from the story and embodying educational-appropriate real-world knowledge.
+Drawing from the results of our human evaluation, the fine-tuned T5-Large exhibits better capabilities in generating QA pairs that suit teachers’ real-world educational expectations of interactive story reading: originating from the story and embodying educational-appropriate real-world knowledge.
 Worth mentioning that T5-Large only consists of $770$ million parameters, whereas Alpaca-7B, Mistral-7B, and Llama 2 in our experiments consist of $7$ billion parameters ($10$ times larger).
-This observation justifies StorySparkQA’s utility in training a task-specific model that caters to parents’ real-world story reading needs on the one hand, and demonstrates the usefulness of combining structured real-world knowledge and free-form narratives in domain-specific tasks such as interactive story reading.
+This observation justifies StorySparkQA’s utility in training a task-specific model that caters to teachers’ real-world story reading needs on the one hand, and demonstrates the usefulness of combining structured real-world knowledge and free-form narratives in domain-specific tasks such as interactive story reading.
 6 Conclusion and Future Work
 In summary, we propose StorySparkQA, an expert-annotated, external-knowledge-enriched QA dataset for children education, by leveraging a novel annotation framework to facilitate scalable expert annotations through structured external knowledge.
-We demonstrate the effectiveness of StorySparkQA through an automated evaluation on various LLMs of generating QA pairs catering to parents’ needs and a human evaluation with children education experts.
-One possible future work is refining the QAG pipeline structures and exploiting LLMs to generate QA pairs that align more closely with parents’ practical needs.
-Another future direction involves using StorySparkQA and language models to develop a human-AI collaborative education system (e.g., an interactive story reading system), aiding parents and educators to formulate personalized questions during story readings, while addressing their language, knowledge, or time constraints.
+We demonstrate the effectiveness of StorySparkQA through an automated evaluation on various LLMs of generating QA pairs catering to teachers’ needs and a human evaluation with children education experts.
+One possible future work is refining the QAG pipeline structures and exploiting LLMs to generate QA pairs that align more closely with teachers’ practical needs.
+Another future direction involves using StorySparkQA and language models to develop a human-AI collaborative education system (e.g., an interactive story reading system) Wang et al.
+(2020, 2019), aiding parents and educators to formulate personalized questions during story readings, while addressing their language, knowledge, or time constraints.
 Also, fine-tuning LLMs (e.g., Llama 3) may lead to better performance on the QAG task, which offers a future direction to refine models’ capabilities in real-world tasks like children’s education.
-7 Acknowledgement
-We thank the Lab of Artificial Intelligence for Education, East China Normal University, Shanghai Institute of Artificial Intelligence for Education, East China Normal University, and Shanghai Science and Technology Innovation Action Plan project (No.
-21511104500) for funding this research.
 7 Limitations
@@
 A.8 LLM Prompts
-To utilize LLMs’ strong reasoning and generation capability as well as control GPT-generated questions as much as possible to meet the needs of parents, we carefully design our prompts.
+To utilize LLMs’ strong reasoning and generation capability as well as control GPT-generated questions as much as possible to meet the needs of teachers, we carefully design our prompts.
 For the QAG pipeline, there are two variations based on the system: (1) Directly generate a QA pair based on a provided story section.